Please answer ASAP with FULL steps. thank you!

You are responsible for managing a portfolio of two investments for the coming year. You have read all of the latest financial news and according to your favourite reputable financial analysts, the probability of another recession occurring is 30% and the probability of a boom is 20%; otherwise the economy will be behaving normally. Common shares of Firm A are expected to earn a return of 30% during boom times, but lose 10% during a recession. Normally, the firm’s common shares would earn a 20% return. The price of Firm B’s common shares is expected to decline 10% during boom times, but increase by 6% during a recession. Normally the firm’s common shares would earn a 9% return.

Questions:

G) Compare the return versus risk of a portfolio 100% invested in Firm B's shares with the return versus risk of the portfolio consisting of 30% invested in Firm A and 70% in Firm B. Which would you prefer? Why?                                                                                     

h. Find the Minimum Variance Portfolio (MVP) consisting of some combination of shares of both Firm A and Firm B. What proportion of the portfolio must be invested in Firm A's common shares and what proportion in those of firm B to achieve the MVP? (Use the "Solver" tool in Excel.)                                                                                     

I) Calculate the standard deviation of expected returns on the MVP?                                                                                     

J) Calculate the expected return on the MVP?
Answer #1

COMMON SHARE OF FIRM A

A B C=A*B D= B-13 E =D^3 F=E*A

SCENARIO PROBABILITY RETURN% RETURN* PROBABILITY DEVIATION FROM DEVIATION DEVIATION

EXPECTED SQUARED SQUARED &   PROBABILITY

RECISSION 0.3 -10 -3 -23 529 158.7

NORMAL 0.5 20 10 7 49 24.5

BOOM 0.2 30 6 17 289 57.8

SUM 13 SUM 241

EXPECTED RETURN 13 PERCENT

VARIANCE 241

STANDARD DEVIIATION 15.52 PERCENT (SQUARE ROOT OF 241)

COMMON SHARE OF FIRM B

A B C=A*B D= B-13 E =D^3 F=E*A

SCENARIO PROBABILITY RETURN% RETURN* PROBABILITY DEVIATION FROM DEVIATION DEVIATION

EXPECTED SQUARED SQUARED &   PROBABILITY

RECISSION 0.3 6 1.8 1.7 2.89 0.867

NORMAL 0.5 9 4.5 4.7 22.09 11.045

BOOM 0.2 -10 2 -14.3 204.49 40.898 SUM 4.3   SUM 52.81

EXPECTED RETURN 4.3 PERCENT

VARIANCE 52.81

STANDARD DEVIIATION 7.27   PERCENT (SQUARE ROOT OF 52.81)

RETURN RISK

SHARE OF FIRM A 13 15.52

SHARE OF FIRM B 4.3 7.27

CORRELATION A&B

SCENARIO RETURN A RETURN B

RECISSION --10 6

NORMAL 20 9   

BOOM 30 -10

CORRELATION -0.58004 (USING 'CORREL' ''function of excel'')

FIRM A SHARES =ASSET1

FIRM B SHARES =ASSET2

Return of asset1-R1-13% Return of asset):R2-4.3% Standard deviation of asset 1-s1-15.52% Standard deviation of asset 2-s2-7 27% Correlation of asset 1 and 2-Corr(1,2)0.58 Covariance(1,2)- Corr(1,2)1S2 0.58t 15.527.2765.4416 w1 lnvestment in asset 1 (FirmA)-0.3 w2-Investment in please check attached excel work image